OpenStudy (mendicant_bias):

Dealing with a conceptual/word problem with Gradients/Vector Fields, I don't understand why the answer is what it is. One moment.

OpenStudy (mendicant_bias):

"Give a formula \[F=M(x,y)i+N(x,y)j\] for the vector field in the plane that has the property that F points toward the origin with magnitude inversely proportional to the square of the distance from (x,y) to the origin. (The field is not defined at (0,0)."

OpenStudy (jhannybean):

Tried drawing it out?

OpenStudy (mendicant_bias):

My confusion is that the answer is \[-a \bigg[x i +yj\bigg](x^2+y^2)^{-3/2}\]

OpenStudy (mendicant_bias):

And I don't understand why it's to the -3/2ths power. I can reasonably imagine you having something like the square of the distance formula in the denominator due to the stated inverse square relationship, but I don't understand why it's to the 3/2ths power. A is just any positive, nonzero arbitrary constant/

ganeshie8 (ganeshie8):

the description of this vector field matches exactly with the gravitational field

OpenStudy (mendicant_bias):

That's dandy, but I don't understand anything more from that, lol.

ganeshie8 (ganeshie8):

but gravity decreases as square of distance in 3D so they are not quite the same

OpenStudy (anonymous):

Try to derive M and N yourself and see if you get the same result when you simplify it.

OpenStudy (mendicant_bias):

I'm really not sure how I should do this "backwards", I just imagined something approximating the formula the way it described it, but I'll take another shot.

OpenStudy (mendicant_bias):

Distance formula in 2D plane:\[\sqrt{x^2+y^2}\] The square of the distance formula is thus\[x^2+y^2\]Having an inverse relationship, you could express the field as\[\frac{1}{x^2+y^2}\]

OpenStudy (mendicant_bias):

(I could elaborate, but it's very clear that there's already something wrong with that, so I'm going to stop there, assuming people immediately have corrections to my thought process.)

OpenStudy (mendicant_bias):

Oh, and let both of those be negative, yeah.

OpenStudy (anonymous):

Lets start with preliminary functions \(M_0, N_0\) which just give a unit vector pointing in the right direction. \(M_0(x, y) = \frac{-x}{|(x, y)|} = \frac{-x}{(x^2 + y^2)^{1/2}} \) \(N_0(x, y) = \frac{-y}{|(x, y)|} = \frac{-y}{(x^2 + y^2)^{1/2}} \) Then we need to add in the factor consisting of its magnitude. Which should be $$\frac{a}{x^2 + y^2}$$ So we have \(M = \frac{-x}{(x^2 + y^2)^{1/2}}\cdot\frac{a}{x^2 + y^2}\) \(N = \frac{-y}{(x^2 + y^2)^{1/2}}\cdot\frac{a}{x^2 + y^2}\)

OpenStudy (mendicant_bias):

Taking a look, one second.

OpenStudy (mendicant_bias):

Alright, yeah, this makes sense. Thanks.

OpenStudy (anonymous):

It always helps to try and derive it yourself. You will be glad you did since you will understand exactly how the formula works in the end.